OpenStudy (anonymous):

graph the equation using the method of your choice. 5x-2y=0

OpenStudy (anonymous):

Solve for y.

OpenStudy (anonymous):

-2y=-5x

OpenStudy (anonymous):

Method of choice? Er. 5x-2y=0 -2y=-5x y=(5/2)x Therefore, the graph is centered on the origin and goes up 5 squares and to the right 2.

OpenStudy (anonymous):

y=(5/2)x

OpenStudy (anonymous):

That's what I got too... thank you